Two sides of a square are on the line x + 10y – 10 = 0 and x + 10y – 5 = 0. Find the area of the square.



1
Expert's answer
2021-11-01T19:25:38-0400

"x+10y-10=0\to y={\frac {10-x} {10}}\to y=-{\frac x {10}}+1"

"x+10y-5=0\to y={\frac {5-x} {10}}\to y=-{\frac x {10}}+{\frac 1 2}"

The picture is presented below



ABCD is the square, then CD is perpendicular to both BC and AD, then it's a distance between them

The area of the ABCD is equal to "CD^{2}"

So, the goal is to find the distance between two given lines (CD).

Since two of the given lines has equal slopes, then they are parallel, and DK equal to "|b{\scriptscriptstyle 1}-b{\scriptscriptstyle 2}|" , where "b{\scriptscriptstyle 1},b{\scriptscriptstyle 2}" - free terms

In the given case "DK=|b{\scriptscriptstyle 1}-b{\scriptscriptstyle 2}|=|1-{\frac 1 2}|=0.5"

"\u2220\u0421DK = \u2220MDA"

"MD=DC"

From two statements above we can conclude that triangles MDA and CDK is equals, so AD = 0.5

From triangle MDA("∠A={\frac \pi 2}"): "AM=AD*tan(\u2220MDA)"

"∠MDA=\pi - ∠MDF \to tan(∠MDA)=tan(\pi -∠MDF)=-tan(∠MDF) = -k ="

"= -(-0.1)=0.1" (k is the slope of the given lines)

"AM=AD*0.1=0.05"

Now from the triangle MDA using the Pythagorian theorem we receive:

"AD^{2}=MA^{2}+AD^{2}=0.2525=CD^{2}" = area of ABCD

The area of the square is 0.2525